• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 22:05:05 on July24,2006

    Entry number 175155

    This entry is a followup to: 175152

    keyword=HRS-L Q2 quench fault (Tech coming in)

    Following up the issue reported in previous halog entry (couldn't get Q2 current to change appropriately), we discovered that we had a communication problem with the HRS-L magnets. The dipole still responded to changing current-set values, *but* we could not enter "P0 Set" values and have them recognized by any of the magnets, *and* we could not manually current-set either Q2 or Q3. So, we rebooted the IOC (IOC 14 @ hareboot10), then all was good again. We regained proper communications via both "P0 Set" and current-sets on all magnets. So then we started normal momentum set procedure by cycling Q2 and Q3 (although now I realize since I am manually setting the fields for Q2 and Q3 at this low momentum, this cycling was not a strictly required procedure). Q3 was all good: went up to 1600 A and brought it back down properly. Q2 however Quenched at just a few hundred Amps. Helium levels appear now to be at 58%...too low.
    Scot is now on his way in to reset Q2 and maybe open the He flow.